# GPTVoiceTasker: LLM-Powered Virtual Assistant for Smartphone

### v1

@@ 5.3. Procedure
We conducted face-to-face user evaluations using an Android device as the experimental tool. On this device, we had the graph of each experimental app populated, which include the majority of app pages and navigation within the app. At the start of the sessions, participants were introduced to all experimental tools via demonstrative videos. The preliminary phase involved practicing basic tasks across all tools, enhancing participants’ familiarity with step-by-step instructions and informative walk-through videos. We also use the searching for exercise tasks in Fig. 1 as the practice tasks, allowing users to achieve this task using each of the tool.
@@ 5.4.1. Overall User Performance
In Fig. 4, we present the average task completion times for each experimental tool. Our GptVoiceTasker stands out with an average completion time of 92.5 seconds, significantly surpassing Voice Access (162.2 seconds) and Voicify (141.9 seconds). This substantial improvement in GptVoiceTasker’s performance can be attributed to two primary factors. Firstly, GptVoiceTasker is better at comprehending user intentions and mapping user commands to the correct actions on specific UI elements, regardless of the command format. In contrast, baseline tools often demand specific command formats, introducing errors in various usages. This issue caused extra time costs as participants needed to seek different ways to express their intentions with the baseline tools. For example, participants tried to tap the option button, in the Notes app with Voice Access by multiple attempts such as “press on the option button”, “press the three-dot icons”, “tap icon for options” before successfully give the right command “tap option”. Secondly, GptVoiceTasker optimizes the performance by automating several steps in one user command, as shown in Table 4. On average, the participants saved 2.2 steps across all six tasks. For instance, in Task 2, GptVoiceTasker efficiently automated the process of searching for Love Yourself song (as in Fig. 5(B)), drawing from a previously stored action designed for searching other songs. This eliminated the need for three steps required for in-app navigation. However, some participants did not realize that they could easily trigger the saved tasks, leading to a missed opportunity for a significant performance boost. In addition, GptVoiceTasker relates to network latency when sending and receiving data from the LLMs API endpoint. This issue could be mitigated with a better network connection.
Figure 5. The comparison between GptVoiceTasker, Voicify, and Voice Access for A) the average cognitive load when using NASA-TLX form (lower is better) *: p < 0.01, **: p < 0.001 and B) Task 2 from the user evaluation with GptVoiceTasker and other baselines.
@@
Figure 5. The comparison between GptVoiceTasker, Voicify, and Voice Access for A) the average cognitive load when using NASA-TLX form (lower is better) *: p < 0.01, **: p < 0.001 and B) Task 2 from the user evaluation with GptVoiceTasker and other baselines.
Figure 6. The comparison between GptVoiceTasker, Voicify, and Voice Access for the System Usability Scale (SUS). *: p < 0.01, **: p < 0.001.
The figure shows the System usability scale result for GPTVoiceTasker and 2 baselines on the scale 1 to 5, where higher is the better for the first five category while lower is better for the last 5 categories. For confidence level score, GPTVoiceTasker achieved 4.28, while Voicify score is 3.5 and Voice Access score is 2.72. For well-integrated score, GPTVoiceTasker achieved 4.44, while Voicify score is 3.17 and Voice Access score is 2.94. For would use frequently score, GPTVoiceTasker achieved 4.17, while Voicify score is 3.44 and Voice Access score is 2.28. For learn to use quickly score, GPTVoiceTasker achieved 4.06, while Voicify score is 3.72 and Voice Access score is 2.67. For easy-to-use score, GPTVoiceTasker achieved 4.28, while Voicify score is 3.06 and Voice Access score is 2.78. For unnecessarily complex score, GPTVoiceTasker achieved 1.67, while Voicify score is 2.83 and Voice Access score is 3.39. For need technical support score, GPTVoiceTasker achieved 1.72, while Voicify score is –2.5 and Voice Access score is 3.72. For inconsistency score, GPTVoiceTasker achieved 1.94, while Voicify score is 3.28 and Voice Access score is 3.83. For cumbersome score, GPTVoiceTasker achieved 1.83, while Voicify score is 3.44 and Voice Access score is 3.89. For need to learn a lot score, GPTVoiceTasker achieved 2.11, while Voicify score is 4 and Voice Access score is 3.94.
Figure 6. The comparison between GptVoiceTasker, Voicify, and Voice Access for the System Usability Scale (SUS). *: p < 0.01, **: p < 0.001.
@@ 5.4.2. Cognitive Load & Usability Ratings
Fig. 5(A) presents an overview of participant feedback regarding their cognitive load levels for each system, assessed using the NASA-TLX form. We conducted a one-way ANOVA statistical analysis, confirming that GptVoiceTasker significantly enhances user performance while reducing frustration levels (p < 0.001). Participants reported decreased mental demand, temporal demand, and effort when using GptVoiceTasker in comparison to the baseline systems. This result signifies a substantial improvement in GptVoiceTasker’s ability to reduce the cognitive load required for operation, aligning with our design goal.
To assess GptVoiceTasker’s usability in comparison to the baseline systems, we employed a one-way ANOVA statistical analysis on collected System Usability Scale (SUS) scores, as depicted in Fig. 6. The analysis verified the enhanced usability of the voice control system, with GptVoiceTasker achieving an average SUS score of 79.861, surpassing Voicify (47.917) and Voice Access (36.528). Participants found GptVoiceTasker easy to use (p < 0.001) and well-integrated (p < 0.001), leading to increased confidence levels (p < 0.001). This remarkable outcome can be attributed to GptVoiceTasker’s ability to effortlessly comprehend natural human commands, reducing the need for extensive training and practice. The lower likelihood of misinterpreting user commands also contributed to the positive results. compared to the baselines, indicating its quick learnability and user satisfaction, thus promoting frequent usage.
@@ 6. Discussion
We introduced GptVoiceTasker as a pioneering example of leveraging LLMs in the development of speech-based virtual assistants. In this section, we delve into the implications and limitations of GptVoiceTasker.
Towards widespread adoption of the voice-centric interface. Advancements in natural language understanding, LLMs like GPT and Bard (Rahaman et al., 2023), have catalyzed the shift towards voice-centric interfaces, extending their use beyond smartphones to wearable devices like smartwatches and AR-VR head-mounted displays. These interfaces not only seamlessly integrate software into daily life but also significantly enhance accessibility for users with visual impairments (Zhong et al., 2014). However, visual-manual methods such as tapping on smartphones or mouse-clicking in desktops have been preferred for their speed and accuracy. Therefore, the transition from the dominant visual-manual interaction to a voice-driven approach presents challenges, stemming not only from the fundamental differences between these interaction modes but also from user unfamiliarity with voice-based interactions. Our tool overcomes this by using LLMs to enhance the intuitiveness of voice interactions. This allows for more intelligent mapping of user intentions to visual elements, easing the shift to voice-assisted interactions and suggesting wider adoption of voice-centric interfaces. Despite the promise, challenges such as the effectiveness of voice recognition in diverse environments still persist. Addressing these will be crucial for the broader adoption of voice-centric interfaces, like smart homes and healthcare. This transition, while challenging, opens new avenues for user interaction and emphasizes the need for continued research in the HCI domain.
@@
LLMs for task automation on user visual interfaces. Research has highlighted the capability of LLMs to provide reasoning based on the UI layout, applying to task automation and testing tools (Feng and Chen, 2023). These models show remarkable capabilities in incorporating extensive knowledge concerning prevalent app design principles and recognizing standard mobile interface elements, including the toolbar, navigation drawer, and bottom navigation bar (Liu et al., 2023b) to significantly enhance proficiency in facilitating precise in-app navigation. Our study highlighted the vital role of spatial information and hierarchical UI representations for LLMs in comprehending semantic connections between diverse UI elements, particularly useful for elements lacking textual information like unlabeled icons or images. In our user study, when tasked with deleting a message lacking a visible delete button, LLM intelligently suggested initiating the process by pressing the unlabelled icon button at the top right, typically the location of the option button, and then selecting “delete” from the ensuing options list. The core of this research lies in the transformation of visual interfaces into textual descriptions that LLMs can process, a critical step for enabling effective task execution based on user inputs. Future research should address the models’ limitations in unconventional UI scenarios and focus on expanding their adaptability across varied interface designs and complex user tasks. Such progress in LLM capabilities is pivotal for advancing user interface automation, leading to more user-friendly and efficient digital experiences.
@@
Towards responsible AI in software systems. In recent years, the remarkable advancements in LLMs have enabled the seamless integration of AI into various software and systems, with the flexibility for fine-tuning and few-shot learning to tackle diverse and challenging downstream tasks. However, this integration raises significant concerns, particularly regarding data privacy and security (Sun et al., 2023). The very nature of AI-integrated systems requires access to data, potentially putting sensitive or confidential information at risk. Put in the context of voice assistants on smartphones, users are sceptical as smartphones contain many personal and sensitive data (Kokolakis, 2017). Tools like GptVoiceTasker can read such on-screen data and further process them to LLMs. To mitigate these risks, several essential measures must be implemented to not only protect users but also build trust, fostering greater adoption of AI-based user interactive systems. First, data de-identification techniques (Liu et al., 2023d) should be applied for on-screen sensitive data, especially for user information such as passwords and personal details. This can be set up prior to usage by the user to hide specific keywords or terms or be detected in real-time using heuristics and deep learning models (Myasnikov and Savchenko, 2019). Furthermore, transparent user consent mechanisms are crucial to informing individuals about data access, usage, and purposes, ensuring they maintain control over their information. Users should have the ability to grant, deny, or adjust permissions as needed to protect their privacy within AI-integrated systems.
@@ 7. Conclusion
In this paper, we introduce GptVoiceTasker, an innovative virtual assistant designed to enhance user interactions and performance on smartphones. GptVoiceTasker leveraged advanced prompt engineering techniques to harness the capabilities of Large Language Models for interpreting user commands and constructing logical reasoning components. GptVoiceTasker further streamlined user interactions by automatically storing previous usages to automate subsequent repetitive tasks. Our experiments demonstrated outstanding command interpretation accuracy and the effectiveness of automated execution based on historical usage. In addition, the user evaluation validated GptVoiceTasker’s high usability in real-world tasks by improving user performance and reducing mental stress load, aligning with our design objectives. As an open-source project, GptVoiceTasker paves the way for future enhancements in virtual assistant intuitiveness, contributing to the evolution of human-computer interactions. Further research includes applying our versatile database execution approach across diverse platforms and operating systems, as well as exploring innovative prompt engineering techniques to fine-tune LLMs for various reasoning tasks.

### v2

@@ 5.3. Procedure
We conducted face-to-face user evaluations using an Android device as the experimental tool. On this device, we had the graph of each experimental app populated, which include the majority of app pages and navigation within the app. At the start of the sessions, participants were introduced to all experimental tools via demonstrative videos. The preliminary phase involved practicing basic tasks across all tools, enhancing participants’ familiarity with step-by-step instructions and informative walk-through videos. We also use the searching for exercise tasks in Figure. 2 as the practice tasks, allowing users to achieve this task using each of the tool.
@@ 5.4.1. Overall User Performance
In Figure. 5, we present the average task completion times for each experimental tool. Our tool stands out with an average completion time of 92.5 seconds, surpassing Voice Access (162.2 seconds) and Voicify (141.9 seconds). This improvement in GptVoiceTasker’s performance can be attributed to two primary factors. Firstly, we can tell that GptVoiceTasker is better at comprehending user intentions and mapping user commands to the correct actions on specific UI elements, regardless of the command format. In contrast, baseline tools often demand specific command formats, introducing errors in various usages. This issue caused extra time costs as participants needed to seek different ways to express their intentions with the baseline tools. For example, participants tried to tap the option button, in the Notes app with Voice Access by multiple attempts such as “press on the option button”, “press the three-dot icons”, “tap icon for options” before successfully give the right command “tap option”. Secondly, GptVoiceTasker optimizes the performance by automating several steps in one user command, as shown in Table 4. On average, the participants saved 2.2 steps across all six tasks. For instance, in Task 2, GptVoiceTasker efficiently automated the process of searching for Love Yourself song (as in Figure. 6(B)), drawing from a previously stored action designed for searching other songs. This eliminated the need for three steps required for in-app navigation. However, some participants did not realize that they could trigger the saved tasks, leading to a missed opportunity for a significant performance boost. In addition, GptVoiceTasker relates to network latency when sending and receiving data from the LLMs API endpoint. This issue could be mitigated with a better network connection.
Figure 6. The comparison between GptVoiceTasker, Voicify, and Voice Access for A) the average cognitive load when using NASA-TLX form (lower is better) *: p ¡ 0.01, **: p ¡ 0.001 and B) Task 2 from the user evaluation with GptVoiceTasker and other baselines.
[…]
@@
Figure 7. The comparison between GptVoiceTasker, Voicify, and Voice Access for the System Usability Scale (SUS). *: p ¡ 0.01, **: p ¡ 0.001.
[…]
The figure shows the System usability scale result for GPTVoiceTasker and 2 baselines on the scale 1 to 5, where higher is the better for the first five category while lower is better for the last 5 categories. For confidence level score, GPTVoiceTasker achieved 4.28, while Voicify score is 3.5 and Voice Access score is 2.72. For well-integrated score, GPTVoiceTasker achieved 4.44, while Voicify score is 3.17 and Voice Access score is 2.94. For would use frequently score, GPTVoiceTasker achieved 4.17, while Voicify score is 3.44 and Voice Access score is 2.28. For learn to use quickly score, GPTVoiceTasker achieved 4.06, while Voicify score is 3.72 and Voice Access score is 2.67. For easy-to-use score, GPTVoiceTasker achieved 4.28, while Voicify score is 3.06 and Voice Access score is 2.78. For unnecessarily complex score, GPTVoiceTasker achieved 1.67, while Voicify score is 2.83 and Voice Access score is 3.39. For need technical support score, GPTVoiceTasker achieved 1.72, while Voicify score is 2.5 and Voice Access score is 3.72. For inconsistency score, GPTVoiceTasker achieved 1.94, while Voicify score is 3.28 and Voice Access score is 3.83. For cumbersome score, GPTVoiceTasker achieved 1.83, while Voicify score is 3.44 and Voice Access score is 3.89. For need to learn a lot score, GPTVoiceTasker achieved 2.11, while Voicify score is 4 and Voice Access score is 3.94.
@@ 5.4.2. Cognitive Load & Usability Ratings
Figure. 6(A) presents an overview of participant feedback regarding their cognitive load levels for each system, assessed using the NASA-TLX form. We conducted a Friedman test (Sheldon et al., 1996) for statistical analysis on the result. Participants reported decreased mental demand, temporal demand, and effort when using GptVoiceTasker in comparison to the baseline systems while achieving better performance. This result show an improvement in GptVoiceTasker’s ability to reduce the cognitive load required for operation, aligning with our design goal.
To assess GptVoiceTasker’s usability in comparison to the baseline systems, we employed Friedman test for statistical analysis on collected System Usability Scale (SUS) scores, as depicted in Figure. 7. The analysis verified the enhanced usability of the voice control system, with GptVoiceTasker achieving an average SUS score of 79.861, surpassing Voicify (47.917) and Voice Access (36.528). Participants found GptVoiceTasker less complex (p ¡ 0.001), less inconsistent (p ¡ 0.001) and well-integrated (p ¡ 0.001), leading to more frequently use (p ¡ 0.001). In addition, participants can learn to use GptVoiceTasker quickly (p ¡ 0.001) as they do not need to learn a lot (p ¡ 0.001). This remarkable outcome can be attributed to GptVoiceTasker’s ability to effortlessly comprehend natural human commands, reducing the need for extensive training and practice. The lower likelihood of misinterpreting user commands also contributed to the positive results.
@@ 6. Discussion
We introduced GptVoiceTasker as an autonomous speech-based virtual assistants. In this section, we delve into the implications and limitations of GptVoiceTasker.
Towards the adoption of the voice-centric interface. The advancements in natural language understanding, particularly through LLMs like GPT and Bard (Rahaman et al., 2023), are propelling the transition towards voice-centric interfaces. These interfaces expand the capabilities smartphones to devices such as smartwatches, AR-VR headsets, and desktops, thereby becoming more integral to everyday activities. While visual-manual methods such as tapping on smartphones or mouse-clicking on desktops are preferred for their speed and accuracy, GptVoiceTasker leverages the power of LLMs to enhance the intuitiveness of voice interactions. This enables more intelligent mapping of user intentions to visual elements, facilitating the shift to voice-assisted interactions and promoting wider adoption of voice-centric interfaces. Additionally, the features introduced by GptVoiceTasker contribute to the domain of voice-centric research on other devices. For example, breaking down one task into a sequence of actions provides more flexibility than fixed intents (e.g., Firefox Voice (Cambre et al., 2021)). The continuous learning from historical usage helps these systems understand user commands better and personalize efficient experiences, such as prefilling patient information for Talk2Care (Yang et al., 2024) and providing personal shortcuts for Firefox Voice (Cambre et al., 2021). Moreover, the implementation of anonymization techniques addresses privacy concerns, particularly for voice-centric interfaces that access the display content from screens.
Voice-centric interfaces also improve accessibility for users with disabilities (Zhong et al., 2014). GptVoiceTasker helps individuals with motor and visual impairments by substituting touch-based interactions with voice commands. For motor impairments, this enables easier task completion on mobile devices without touching the screen. Additionally, GptVoiceTasker provides voice shortcuts for visually impaired users, allowing quicker navigation of familiar screens. For example, users can find specific buttons with commands instead of clicking through each one with Talkback. Natural interaction methods could also benefit individuals who struggle with technology and elderly users.
Despite the promise, challenges such as the effectiveness of voice recognition in diverse environments still persist. Addressing these will be crucial for the broader adoption of voice-centric interfaces, like smart homes and healthcare. This transition, while challenging, opens new avenues for user interaction and emphasizes the need for continued research in the HCI domain.
@@
LLMs for task automation on user visual interfaces. Research has highlighted the capability of LLMs to provide reasoning based on the UI layout, applying to task automation and testing tools (Feng and Chen, 2023). These models show remarkable capabilities in incorporating extensive knowledge concerning prevalent app design principles and recognizing standard mobile interface elements, including the toolbar, navigation drawer, and bottom navigation bar (Liu et al., 2023b) to enhance proficiency in facilitating precise in-app navigation. Our study highlighted the vital role of spatial information and hierarchical UI representations for LLMs in comprehending semantic connections between diverse UI elements, particularly useful for elements lacking textual information like unlabeled icons or images. In our user study, when tasked with deleting a message lacking a visible delete button, LLM intelligently suggested initiating the process by pressing the unlabelled icon button at the top right, typically the location of the option button, and then selecting “delete” from the ensuing options list. The core of this research lies in the transformation of visual interfaces into textual descriptions that LLMs can process, a critical step for enabling effective task execution based on user inputs. Future research should address the models’ limitations in unconventional UI scenarios and focus on expanding their adaptability across varied interface designs and complex user tasks. As models grow, visual models (e.g., GPT-4v) can process images to further enhance the accuracy of such interactions. However, optimizing the usage of visual models to balance accuracy and efficiency is crucial to compensate for the drop in response time, thereby improving their practical application in real-time scenarios. Such progress in LLM capabilities is pivotal for advancing user interface automation, leading to more user-friendly and efficient digital experiences.
@@
Towards responsible AI in software systems. In recent years, the remarkable advancements in LLMs have enabled the seamless integration of AI into various software and systems. However, this integration raises significant concerns, particularly regarding data privacy and security (Sun et al., 2023). The very nature of AI-integrated systems requires access to data, potentially putting sensitive or confidential information at risk. Put in the context of voice assistants on smartphones, users are sceptical as smartphones contain many personal and sensitive data (Kokolakis, 2017). Tools like GptVoiceTasker can read such on-screen data and further process them to LLMs. To mitigate these risks, it is essential to implement several key measures, not only to protect users but also to build trust, thereby fostering greater adoption of AI-based interactive systems. GptVoiceTasker represents a pioneering effort in voice-assistive research by applying personal information anonymization to protect user privacy when using LLMs for logical tasks. Additionally, when executing actions on behalf of users, voice assistants must operate responsibly, ensuring that actions do not adversely affect users. This involves seeking explicit user confirmation for decisions, particularly in scenarios where actions may have significant implications, such as replying to important emails or transferring money. Future work in this field should focus on identifying sensitive actions and prompting user confirmation while maintaining a seamless user experience.
@@ 7. Conclusion
In this paper, we introduce GptVoiceTasker, an innovative virtual assistant designed to enhance user interactions and performance on smartphones. GptVoiceTasker leveraged advanced prompt engineering techniques to harness the capabilities of LLMs for interpreting user commands and constructing logical reasoning components. GptVoiceTasker further streamlined user interactions by automatically storing previous usages to automate subsequent repetitive tasks. Our experiments demonstrated outstanding command interpretation accuracy and the effectiveness of automated execution based on historical usage. In addition, the user evaluation validated GptVoiceTasker’s high usability in real-world tasks by improving user performance and reducing mental stress load, aligning with our design objectives. As an open-source project, GptVoiceTasker paves the way for future enhancements in virtual assistant intuitiveness, contributing to the evolution of human-computer interactions. Further research includes applying our versatile database execution approach across diverse platforms and operating systems, as well as exploring innovative prompt engineering techniques to fine-tune LLMs for various reasoning tasks.